Grade 9Magical Mathematics[Interesting Approach]

relation between height , curved surface area and volume of cone

Abijeeth Thirumalainathan
13 Years agoGrade 9

2 Answers

Chetan Mandayam Nayakar
13 Years ago

volume=(1/3)(pi)(radius of base)2(height)

curved surface area=(pi)(r)(l), r is radius of base and l is length of straight line connecting apex of cone with point on edge of base

lokesh soni
13 Years ago

for cone

curved surface area =pi*r*√(r2+h2)

volume= 1/3(pi*r2h)